### v1

# Language-Specific Neurons: The Key to Multilingual Capabilities in Large Language Models

Abstract
Large language models (LLMs) demonstrate remarkable multilingual capabilities without being pre-trained on specially curated multilingual parallel corpora. It remains a challenging problem to explain the underlying mechanisms by which LLMs process multilingual texts. In this paper, we delve into the composition of Transformer architectures in LLMs to pinpoint language-specific regions. Specially, we propose a novel detection method, language activation probability entropy (LAPE), to identify language-specific neurons within LLMs. Based on LAPE, we conduct comprehensive experiments on two representative LLMs, namely LLaMA-2 and BLOOM. Our findings indicate that LLMs’ proficiency in processing a particular language is predominantly due to a small subset of neurons, primarily situated in the models’ top and bottom layers. Furthermore, we showcase the feasibility to “steer” the output language of LLMs by selectively activating or deactivating language-specific neurons. Our research provides important evidence to the understanding and exploration of the multilingual capabilities of LLMs.
Language-Specific Neurons:
The Key to Multilingual Capabilities in Large Language Models
Tianyi Tang1,3 * †thanks: * This work was done during internship at MSRA., Wenyang Luo1 †thanks: † Equal contribution., Haoyang Huang2, Dongdong Zhang2 Xiaolei Wang1, Wayne Xin Zhao1,3 🖂thanks: 🖂 Corresponding author , Furu Wei2, Ji-Rong Wen1,3 1 Gaoling School of Artificial Intelligence, Renmin University of China 2 Microsoft Research Asia, China 3 Beijing Key Laboratory of Big Data Management and Analysis Methods steventianyitang@outlook.com  luowenyang@ruc.edu.cn  wxl1999@foxmail.com {haohua,dozhang,fuwei}@microsoft.com    batmanfly@gmail.com
1 Introduction
The brain has its own language for testing the structure and consistency of the world.
Carl Sagan
The pursuit of multilingual capabilities, mirroring our world’s linguistic diversity, is a critical research objective that paves the way for information democratization across linguistic divides. The emergence of pre-trained language models (PLMs) such as mBERT Devlin et al. (2019) and XLM-R Conneau et al. (2020a) has marked a significant shift towards enhanced multilingual understanding. Furthermore, large language models (LLMs), such as GPT-4 Achiam et al. (2023) and PaLM-2 Anil et al. (2023), have recently demonstrated more excellent multilingual capabilities in language understanding, reasoning, and generation, despite being predominantly trained in English corpora.
Figure 1: An illustration of region distribution of activated neurons when predicting the next word in language models across different languages. Here, colored circles denote activated neurons. When Chinese-specific neurons are deactivated (denoted by $\otimes$), the model may produce outputs in English.
Existing studies Pires et al. (2019); Conneau et al. (2020b) have mainly explored how multilingual PLMs (e.g., mBERT) possess semantic alignment capabilities across languages despite the absence of multilingual parallel corpora. They have identified several critical factors that influence cross-lingual transfer, including training data (e.g., overlapped tokens) and training settings (e.g., shared parameters) Dufter and Schütze (2020); Philippy et al. (2023). Nevertheless, the underlying mechanisms by which the model itself process diverse languages at the composition level continue to be an area of vigorous investigation.
To develop a deeper understanding of the multilingual capabilities of LLMs, we draw inspiration from the neurobiological underpinnings of human language faculties Friederici (2011); Parr et al. (2022); Khanna et al. (2024). Specific regions within the human brain, such as Broca’s area and Wernicke’s area have been identified to support particular language functions. To make an analogy with human’s language functions, we posit that regions within the language models can be delineated into two primary components: language-agnostic regions that encompass universal knowledge and pragmatics principles, and language-specific regions that handle language-specific vocabulary, grammar, and idiomatic expressions. Figure 1 presents such a conceptual illustration of region distribution in LLMs posited by us. Actually, language-agnostic regions have been widely explored in existing literature, including knowledge storing Dai et al. (2022) and task handling Wang et al. (2022). However, language-specific regions, especially those supporting multilingual capacities, have been seldom studied in existing literature of of LLMs, which is the focus of our research.
In this work, we first propose a novel detection method called language activation probability entropy (LAPE) to identify language-specific neurons within LLMs. This method involves computing the activation likelihood of individual neurons in response to corpora across different languages. Subsequently, we select neurons with lower language activation probability entropy as language-specific neurons, i.e., those having a higher activation probability for one or two particular languages and a lower probability for others. Furthermore, based on the proposed LAPE method, we have conducted a systematic study with language-specific regions of two popular open-sourced LLMs, leading to several major findings:
$\bullet$ First, the proficiency of an LLM in processing a particular language can be significantly impacted by a minuscule proportion of its neurons. Deactivating such language-specific neurons leads to a remarkable degradation in the model’s understanding and generation abilities for that language.
$\bullet$ Second, neurons specific to individual languages are predominantly located in the bottom and top layers of LLMs. The bottom layers mainly serve to process the input from various languages into the unified semantic space of a high-resource language (e.g., English), while the top layers project the semantic content (after the processing of middle layers) into the respective tokens in the corresponding vocabulary of each language.
$\bullet$ Third, we demonstrate the potential to “steer” the output language of LLMs by selectively activating and/or deactivating certain neurons. Our approach could provide a promising solution to mitigate the off-target issue (e.g., the tendency of LLaMA-2 to reply in English to Chinese queries), while stimulating the capabilities of cross-lingual generation tasks.
To the best of our knowledge, it is the first study that investigates language-specific regions inside LLMs and analyzes the how these regions influence LLMs’ capabilities to process multilingual texts. We introduce the concept of “language-specific neurons” and propose language activation probability entropy to identify such neurons in LLMs. We make available the identified language-specific neurons and corresponding code at https://github.com/RUCAIBox/Language-Specific-Neurons.
2 Identifying Language-Specific Regions
2.1 Background
Currently, LLMs are predominantly developed on an auto-regressive Transformer architecture Vaswani et al. (2017), in which the basic building blocks are the multi-head self-attention (MHA) and the feed-forward network (FFN). Given the hidden state $\bm{h}^{i-1}\in\mathbb{R}^{d}$ of ($i-1$)-th layer of a specific token, the MHA module inside the $i$-th layer can be expressed as follows:
$\tilde{\bm{h}}^{i}=\text{Attn}(\bm{h}^{i-1}\bm{W}^{i}_{q},\bm{H}^{i-1}\bm{W}^{i}_{k},\bm{H}^{i-1}\bm{W}^{i}_{v})\cdot\bm{W}^{i}_{o},$ (1)
where $\bm{W}^{i}_{q}$, $\bm{W}^{i}_{k}$, $\bm{W}^{i}_{v}$, and $\bm{W}^{i}_{o}$ represent the trainable parameters, and $\bm{H}^{i-1}$ stands for the hidden states in the previous layer of the whole sequence. Subsequently, the FFN module is described by the following formulation:
$\bm{h}^{i}=\text{act\_fn}(\tilde{\bm{h}}^{i}\bm{W}^{i}_{1})\cdot\bm{W}^{i}_{2},$ (2)
[…]
$\bm{h}^{i}=(\text{act\_fn}(\tilde{\bm{h}}^{i}\bm{W}^{i}_{1})\otimes\tilde{\bm{h}}^{i}\bm{W}^{i}_{3})\cdot\bm{W}^{i}_{2}.$ (3)
[…]
2.2 Language Activation Probability Entropy
In existing research, neurons within the FFN modules are found to be capable of storing factual knowledge Dai et al. (2022), encoding positional information Voita et al. (2023), responding to particular syntactic triggers Gurnee et al. (2024), etc. Inspired by these findings, we posit that there exist specific neurons in LLMs for multilingual processing. Next, we introduce a new detection method based on language activation probability entropy (LAPE) to identify language-specific neurons.
Our research primarily focuses on pre-trained base models (e.g., LLaMA-2 and BLOOM), rather than fine-tuned models that have undergone instruction tuning or RLHF (e.g., LLaMA-2-Chat and BLOOMZ), which helps reduce other influencing factors. Specially, we feed existing LLMs with multilingual texts, each written in a single language. For the $j$-th neuron in the $i$-th layer, we then compute the activation probability when processing texts in language $k$:
$p^{k}_{i,j}=\mathbb{E}\left(\mathbb{I}(\text{act\_fn}(\hat{\bm{h}}^{i}\bm{W}^{i}_{1})_{j}>0)\mid\text{language }k\right),$ (4)
where $\mathbb{I}$ is the indicator function. The activation probability is empirically estimated by the likelihood that the neuron’s activation value exceeds zero. Subsequently, we can obtain the distribution $\bm{p}_{i,j}=(p^{1}_{i,j},\dots,p^{k}_{i,j},\dots,p^{l}_{i,j})$ for each neuron, indicating its probability of activation for each language. To convert $\bm{p}_{i,j}$ into a valid probability distribution, we apply L1 normalization, yielding $\bm{p}_{i,j}^{\prime}$. The entropy of this distribution, which we refer to as language activation probability entropy, is computed to quantify the neuron’s language activation reaction:
$\text{LAPE}_{i,j}=-\sum_{k=1}^{l}p^{\prime k}_{i,j}\log(p^{\prime k}_{i,j}).$ (5)
We designate neurons with low LAPE scores as “language-specific neurons”, as they demonstrate a predilection for activation in response to one or two languages, while showing reduced activation probabilities for others.
In implementation, we collect multilingual corpora sourced from Wikipedia, a widely recognized and high-quality resource for diverse languages, and sample documents to create a dataset comprising 100 million tokens for each language. Subsequently, we input these tokens into a target LLM and follow Equations 4 and 5 to compute the LAPE score for individual neurons. Finally, we select neurons that fall within the lowest percentile of LAPE scores, specifically targeting the bottom 1%. To refine our selection, we further impose a predefined threshold to exclude neurons exhibiting negligible activation probability: a neuron is deemed specific to language $k$ if its corresponding activation probability $p^{k}_{i,j}$ surpasses the threshold.
3 Experiments
In this section, we present empirical evaluation to substantiate the efficacy of our proposed LAPE method and elucidate the impact of language-specific neurons on multilingual capacities.
3.1 Experimental Setup
Models.
We conduct the study based on two public LLMs: LLaMA-2 Touvron et al. (2023b), primarily pre-trained on English texts and known for its excellence as a language foundational model, and BLOOM Scao et al. (2022), known for its multilingual proficiency with a more balanced language distribution. We investigate multiple versions of LLaMA-2, specifically the 7B, 13B, and 70B models, containing approximately 352K, 553K, and 2.29M neurons, respectively. For BLOOM, we select the 7.1B version, consisting of roughly 492K neurons. The languages selected include English (en), Simplified Chinese (zh), French (fr), Spanish (es), Vietnamese (vi), Indonesian (id), and Japanese (ja). We exclude ja for BLOOM since it has not been pre-trained on Japanese copora.
Dataset.
Our analysis of language-specific neurons is conducted across two distinct dimensions:
$\bullet$ Language modeling: We assess the multilingual language modeling capability using perplexity (PPL) scores on Wikipedia corpora. Our dataset comprises one million tokens per language, all sourced after September 2022 to ensure the content has not been included in the training sets of either LLaMA-2 or BLOOM.
$\bullet$ Open-ended generation: To evaluate model’s multilingual generation capabilities in real-world scenarios, we translate a set of questions from the Vicuna dataset Chiang et al. (2023) into the target languages. These questions span a broad spectrum of topics, excluding mathematics and coding queries to maintain a focus on language processing proficiency, avoiding confounding variables. We utilize greedy search with repetition penalty of 1.1 to generate output and then employ GPT-4 to assess the output text on a scale ranging from 1 to 10 following Zheng et al. (2023).
(a) LAPE
(b) LAVE
(c) PV
(d) RS
Figure 2: Impact of four identification methods on the PPL increase of LLaMA-2 (7B). The element at the $i$-th row and $j$-th column is the PPL change for language $j$ due to perturbations in a specific region of language $i$.
Identification methods.
For comparison, we consider the following methods for identifying language-specific regions:
(a) Language activation probability entropy (LAPE, ours): The pertinent details are provided in Section 2.2. The threshold is set to the 95-th percentile of all activation probabilities. For instance, in the case of LLaMA-2 (70B), threshold is established at 0.515. This stipulates that the neurons we select are required to exhibit an activation probability exceeding 0.515 for at least one language.
(b) Language activation value entropy (LAVE): This is a variant of our proposed method, wherein we substitute the activation probability with the mean activation value across languages. Similarly, we normalize the activation value and calculate the entropy to find neurons with high activation value in response to particular languages.
(c) Parameter variation (PV): By extending the work of Zhao et al. (2023a), we compare the model parameters before and after monolingual instruction tuning to identify language-specific parameters. In particular, we train individual models on the Alpaca instruction datasets Taori et al. (2023) and its multilingual version Chen et al. (2023b) which comprise 52,000 instances for each target. These models undergo training for two epochs, with a batch size of 128 and a constant learning rate of 1e-5. We mainly consider the parameters inside the MHA and FFN modules, i.e., the weight matrices in Equations 12, and 3. We compute the rate of change across various languages, and select parameters that exhibit a low rate of change in one or two languages but a high rate in others. In detail, we refine the change ratio by subtracting the maximum value and then conduct L1 normalization for entropy calculation.
(d) Random selection (RS): Additionally, we add a baseline to randomly select neurons for each language, serving as a reference for different methods.
3.2 Main Perturbation Experiments
(a) LLaMA-2 (7B)
(b) LLaMA-2 (13B)
(c) LLaMA-2 (70B)
(d) BLOOM (7B)
Figure 3: Applying our LAPE method to different model types and sizes.
In this part, we conduct the perturbation experiments by deactivating the identified language-specific regions. Specially, for all the comparison methods in Section 3.1, we identify 1% of the neurons or parameters and treat them as language-specific regions. We then set the activation values of these neurons to zero, or directly zero the parameters to assess the impact on the model’s multilingual capacities based on language modeling and open-ended text generation tasks.
Figure 2 presents the perturbation results (measured by PPL change on language modeling task) of different identification methods on LLaMA-2. Overall, we can see that large impact values for LAPE mainly occur in diagonal entries. It indicates that our LAPE method is adept at identifying language-specific neurons. Deactivating neurons associated with a specific language predominantly affects the PPL results of that language, with negligible effects on others. In contrast, the variant utilizing activation values (LAVE) causes cross-lingual interference (e.g., the entry $\langle id,es\rangle$), and the rest two methods PV and RS fails to exhibit clear language-specific patterns.
zh fr es vi id ja
normal 4.30 4.19 3.51 3.70 4.16 2.86
zh 2.46 3.56 2.96 3.64 3.56 2.31
fr 3.69 2.50 2.29 3.01 3.59 2.76
es 3.51 2.57 2.01 3.14 3.34 2.56
vi 3.93 3.19 2.49 2.74 3.59 2.74
id 3.67 3.10 2.67 3.21 2.84 2.80
ja 3.21 3.69 3.07 3.49 3.37 1.84
Table 1: Performance of LLaMA-2 (70B) on the multilingual Vicuna as evaluated by GPT-4. The “normal” row is baseline scores without deactivation. Subsequent rows are scores with deactivation of specific neurons.
We further investigate whether our method is effective across different model sizes and model types. As depicted in Figure 3, the language-specific degradation patterns remain evident for three LLaMA models (7B, 13B, 70B) and a BLOOM model (7B). An interesting find is that neurons in larger models tend to be specialized for a single language rather than being shared among two or more languages. Furthermore, we can find that there exists a high correlation between Chinese and Japanese: when we deactivating the neurons specific to one language, the performance of the other language would be affected (e.g., the entries $\langle zh,ja\rangle$ and $\langle ja,zh\rangle$). By inspecting into the identified neurons of the two languages, we note that a substantial amount of neurons (approximately 25%) actually overlap for both languages. It is likely because Chinese and Japanese partially share common characters. In addition, it can be observed that our LAPE method leads to similar findings on BLOOM-7B (Figure 3(d)). These results further verify the effectiveness of generality of our proposed LAPE method.
When employing LAPE to open-ended generation tasks based on the Vicuna dataset, we can clearly observe from Table 1 that deactivating language-specific neurons significantly impairs the generation capabilities in the target language. We further provide an illustrative example in Table 2 of the model’s response to a question in Simplified Chinese when the neurons associated with Simplified Chinese are deactivated. We can see a chaotic mixture of Traditional Chinese characters and redundant English phrases, indicating a severe degradation of language capacity in Simplified Chinese.
Question
你是一位登上珠穆朗玛峰顶峰的登山者。描述一下你的情绪和从高处看到的景色。
(Translation: You are a mountain climber reaching the summit of Mount Everest. Describe your emotions and the view.)
Normal output
我是一个登上珠穆朗玛峰顶峰的登山者。当我站在山顶时,我感到非常兴奋和自豪。…
(Translation: I am a climber who has reached the summit of Mount Everest. When I stood on the top of the mountain, I felt very excited and proud. …)
Deactivated output
我是一個登上珠穆朗瑪峰頂峰的登山者。 I am a mountaineer who has climbed to the top of Mount Everest. 當我站在珠my朗ma峰頂峰,我感到非常興奮和欣慰。 …
Table 2: Illustration of LLaMA-2-70B responses to a question in Simplified Chinese. The text in black is model’s actual output and text in gray is our self-added translation. The deactivated output is the generation when neurons of Simplified Chinese are deactivated.
3.3 Further Analysis
After presenting the main experiments, we further conduct detailed analysis experiments to investigate language-specific neurons and their impact. Unless specified, all analysis results are obtained based on LLaMA-2 (70B).
3.3.1 Distribution and Identification Ratio
en zh fr es vi id ja
836 5,153 6,082 6,154 4,980 6,106 5,216
Table 3: The number of neurons in each language.
Neuron distribution across languages.
After running our LAPE method on LLaMA-2 (70B), we identify around 23K language-specific neurons. The distribution of these neurons across individual languages is detailed in Table 3. Since neurons may be shared by multiple languages, the sum of language-specific neurons actually surpass 23K. Overall, except English, the distribution of neurons is relatively even across languages. However, the number of English-specific neurons is much smaller than the other languages. We speculate that English is the dominant language in LLaMA-2, and thus it requires fewer neurons to support the specific language ability.
Figure 4: Change in PPL across different languages upon incremental number of language specific neurons when deactivating French neurons.
Increasing threshold ratios for identification.
In Section 3.2, we consider a mere 1% of the neurons as being language specific. We further vary the selection ratio of language-specific neurons from 0 to 10%, to examine its impact on multilingual processing. Here, we select French for study, while the results on the other languages are similar. When deactivating neurons specific to French, we observe a significant increase in the PPL on French in Figure 4, while the impact on the rest languages are relatively limited with the exception of Spanish. it is consistent with our intuition: the performance of the being perturbed language and its related (or similar) language would be severely harmed.
3.3.2 Structural Distribution Analysis
In this part, we analyze how language-specific neurons are distributed across different layers.
Language processing is concentrated at bottom and top layers.
In Figure 5, we present the layer-wise distribution of language-specific neurons across various layers, which has a skewed “U” shape. This finding indicates that language-specific neurons have a pronounced concentration in the top and bottom layers. Specifically, the second layer is composed of approximately 7,000 neurons, while layers 5 through 47 only contain about 100 neurons each. Further, the neuron count gradually increases, with the final four layers each comprising over 1,000 neurons. The complete statistics about the layer-wise distribution across various languages are reported in Table 10 of the Appendix.
Figure 5: Distribution of language-specific neurons across different layers in LLaMA-2 (70B).
Why such a skewed distribution?
To understand why such a skewed distribution occurs, we seek explanations from multilingual semantic representation by exploring how multilingual aligned texts are represented across the layers. Specially, we employ the multilingual Vicuna dataset (Section 3.1), comprising of aligned texts in different languages. Given a group of aligned texts, we feed them into the LLM and obtain the sentence embeddings for each layer. We then compute the mean sentence embedding similarity (SES) among the aligned texts across language in Figure 6. Interestingly, the SES curve shows an opposite trend with the distribution of language-specific neurons. At the beginning, the similarity quickly increases, then reaches the peak, and gradually decreases to a small value. This finding shows: at the bottom layers, the LLM needs to map aligned texts of different languages into the shared representation space, thus requiring more language-specific neurons for semantic transformation; while at top layers serving for token generation, the LLM needs to handle vocabulary mapping, which also requires the support of more language-specific neurons.
Figure 6: The mean SES between all language pairs and total language neuron numbers across layers.
3.3.3 Language Dominance Analysis
Since the high-resource language (i.e., English) in the LLaMA-2 training corpus has a surprisingly smaller number of neurons than other languages, as indicated in Table 3, we speculate that there might exist some dominance relations between high-resource and low-resource languages, which depends on the composition of pre-training corpus.
Language dominance measurement.
Inspired by Xu et al. (2023), we transfer the sentence embeddings across different languages into the same space around a target language, and examine how texts from the other languages are aligned to the texts of the target language. Firstly, for each language $k$, we compute the mean sentence embeddings of all its texts, and obtain $\bm{v}_{k}^{i}$ as the language vector of $k$ at $i$-th layer. Then we follow the same formula proposed by Xu et al. (2023) to conduct the space mapping for each text from language $k$:
$\hat{\bm{h}}_{k}^{i}=\bm{h}_{k}^{i}-\bm{v}_{k}^{i}+\bm{v}_{c}^{i},$ (6)
where $\hat{\bm{h}}_{k}^{i}$ denotes a transformed embedding of some text in the $i$-th layer of language $k$. Here, we specify $c$ as the target language, and compute the mean sentence embedding similarity (SES, Section 3.3.2) over all the sentence pairs between languages $k$ and $c$, based on the transformed representations in Eq. 6. A larger SES score indicates language $c$ has a larger dominance degree.
(a) LLaMA-2 (70B)
(b) BLOOM (170B)
Figure 7: Dominance scores (mean SES) across layers when different languages serve as the target language.
Low-resource languages are centered around high-resource languages.
To compute the dominance degree, we still use the multilingual Vicuna dataset (Section 3.1). From the results of LLaMA-2 (70B) in Figure 6(a), we can see that the mean SES is obviously higher than all other languages when the target language is English. As English is the high-resource language of LLaMA-2, low-resource languages need to be aligned with English for achieving better performance. When it comes to BLOOM (170B) in Figure 6(b), several languages (e.g., English and Chinese) shows dominance, since it is originally pre-trained on multilingual corpora.
Metrics Settings zh fr es vi id ja
Language normal 0.87 0.73 0.81 0.60 0.40 0.79
accuracy steered 0.99 0.90 0.93 0.97 0.99 1.00
Content normal 4.30 4.19 3.51 3.70 4.16 2.86
quality steered 4.57 4.35 4.02 3.57 4.28 2.91
Table 4: The language accuracy and content score of the normal output and the steered output by activating language-specific neurons. Language accuracy is computed by whether the model responds in a given language using the langdetect package and the content quality is measured by GPT-4.
3.3.4 Case Study
Finally, we explore the possibility to “steer” the output language of LLMs to mitigate the off-target problem and facilitate cross-lingual generation.
Researchers have observed that when prompting in one language, language models may generate responses in a different language, such a phenomenon referred to as the off-target language issue Gu et al. (2019); Sennrich et al. (2023). We speculate that language-specific neuron might not be activated in this case. Thus, we attempt to manually activating these neurons by increasing their activation value to the average activation value of this language. As shown in Table 4, it is evident that the model demonstrates a significantly higher propensity to respond in the expected language, as indicated by the language accuracy. Additionally, the generation quality, as assessed by GPT-4 metrics, shows remarkable improvement.
Moreover, we attempt to steer cross-lingual generation by posing a question in one language and answering it in another. As presented in Table 5, we deactivate Spanish neurons and activate Chinese neurons when replying to a question in Spanish. The model successfully generates fluent Chinese text, showing the feasibility of altering the output language through neuron-level manipulation.
Question
?‘Cómo puedo mejorar mis habilidades de gestión del tiempo?
(Translation: How can I improve my time management skills?)
Steered output in Chinese
Time管理是一个非常重要的技能,因为它可以提高你的生产力并使你更有效地完成任务。下面是一些提示来改进你的时间管理技能:1、制定计划:每天开始时制定清单并将任务分配给特定时间段。这样可以确保你在每天都有足够的时间来完成所有任务。2、设置目标:设置明确而可实现的目标可以帮助你保持专注并提高效果。3、抵制分心:当你正在工作时,请不要被其他事情分心。4、利用技术:……
(Translation: Time management is a very important skill because it increases your productivity and allows you to complete tasks more efficiently. Here are some tips to improve your time management skills: 1. Make a plan: Make a list at the beginning of each day and assign tasks to specific time periods. This ensures that you have enough time each day to complete all your tasks. 2. Set goals: Setting clear and achievable goals can help you stay focused and improve your results. 3. Resist distractions: When you are working, please don’t be distracted by other things. 4. Utilize technology: …)
Table 5: An illustrative example of asking question in Spanish and answering it by deactivating Spanish neurons and activating Chinese neurons.
4 Related Work
Multilingual ability.
Multilingual ability has been widely studied since the emergence of pre-trained language models Devlin et al. (2019); Xue et al. (2021). Recently, LLMs Zhao et al. (2023b); Nguyen et al. (2023) showcase more excellent multilingual capabilities even without explicit language alignment Kulshreshtha et al. (2020); Cao et al. (2020). Extensive research has been made to investigate the factors that influence models’ multilingual ability Philippy et al. (2023). For example, linguistic similarity has been examined Pires et al. (2019); Dufter and Schütze (2020); Wang et al. (2019); Conneau et al. (2020b), which is generally believed to correlate with models’ cross-lingual ability. Specially, “word order” shows some contradictions about whether it really affects multilingual ability Pires et al. (2019); Deshpande et al. (2022). Not only limited to language property, training settings have also been considered Lauscher et al. (2020); Ahuja et al. (2022).
Existing work has explored language-agnostic (or language-shared) components within multilingual models. For example, researchers concentrate on shared knowledge across various languages Stanczak et al. (2022); Chen et al. (2023a); Zhao et al. (2023a); Bricken et al. (2023). However, the exploration of language-specific components within LLMs remains an under-investigated area.
Neuron analysis.
Neuron analysis has gained much attention in recent years following neurobiological human brain Friederici (2011); Parr et al. (2022). Originating from vision models Bau et al. (2020); Mu and Andreas (2020), neuron analysis views the neuron activation as the recall of learned knowledge Sajjad et al. (2022). Researchers widely adopt them to analyze the source of some specific abilities or skills language models, including sentiment analysis (Radford et al., 2017), knowledge storing (Dai et al., 2022), task solving Wang et al. (2022). Further, some recent studies have also found that certain neurons can convey special contexts Gurnee et al. (2023); Bills et al. (2023), such as positions Voita et al. (2023) and linguistic related properties Bau et al. (2018); Xin et al. (2019); Dalvi et al. (2019, 2020). Moreover, Gurnee et al. (2024) use the Pearson correlation to calculate neuron similarity and find some universal neurons across models. In contrast, neuron analysis is not widely utilized in multilingual areas, and we extend this method to LLMs and unveil the mechanism of their multilingual ability.
5 Conclusion
Despite the impressive multilingual capabilities demonstrated by LLMs, the understanding of how these abilities develop and function remains nascent. In this paper, we introduced a novel detection method, i.e., language activation probability entropy (LAPE), to pinpoint language-specific neurons within LLMs. LAPE assesses the response of individual neurons to various languages, selecting those with a propensity for activation when exposed to one or two languages. Based on LAPE, we further conducted extensive experiments to investigate the multilingual capabilities of LLMs. Specially, we have found that an LLM’s proficiency in processing different languages is significantly influenced by a small subset of neurons, which are mainly located in the model’s top and bottom layers. We have further demonstrated that the output language of LLMs can be directed by selectively enabling or disabling these language-specific neurons. For future work, we aim to leverage these findings to enhance knowledge transfer between major and minor languages and devise efficient strategies for continual pre-training to better accommodate specific languages.
Limitations
In this study, we employ language activation probability entropy as a metric to identify language-specific neurons. However, it is important to note that our method is relative to the presence of multiple languages. In scenarios where only a single language is present, establishing an absolute threshold to determine the language-relatedness of neurons is not feasible. Moreover, the criteria for distinguishing between high-resource and low-resource languages within the model warrant further investigation. The model’s possibility to managing a large number of languages, as well as the differences between various languages, represents a promising avenue for future research. Finally, our research has only begun to explore the possibility for directing the output language of the model. Developing strategies to harness these identified neurons for enhancing the model’s multilingual proficiency is still worth exploring.
Appendix A Appendix
Table 7 compiles the statistics of pre-training corpora in LLaMA-2 Touvron et al. (2023b) and BLOOM Scao et al. (2022).
We list the number of language-specific neurons across different layers of BLOOM (7B), LLaMA-2 (7B), LLaMA-2 (13B), and LLaMA-2 (70B) in Tables 6, 8, 9, and 10.
#Layer en zh fr es vi id
1 0 0 0 0 0 0
2 0 0 0 0 0 0
3 0 0 0 0 0 0
4 0 0 0 0 0 0
5 0 1 0 0 0 0
6 0 0 0 0 0 0
7 0 1 0 0 1 0
8 0 0 0 0 1 0
9 0 0 1 0 0 0
10 1 2 1 1 0 0
11 0 0 2 2 1 1
12 0 0 0 0 0 0
13 1 1 1 1 0 0
14 0 1 1 1 0 0
15 1 2 1 0 0 0
16 1 0 2 2 1 2
17 3 0 1 1 0 0
18 2 1 3 2 0 1
19 3 4 3 4 2 2
20 1 1 2 1 1 1
21 11 7 7 8 3 4
22 8 8 9 9 7 9
23 9 19 11 12 12 15
24 21 24 24 24 26 46
25 24 34 24 28 35 90
26 24 37 47 54 40 180
27 34 46 66 93 70 330
28 62 79 106 151 83 562
29 86 126 155 240 103 817
30 153 259 213 284 165 763
Table 6: Neuron number per layer of BLOOM (7B).
Language Code Family BLOOM Ratio LLaMA-2 Ratio
English en Indo-European 33.68% 89.70%
Chinese zh Sino-Tibetan 18.13% 0.13%
French fr Indo-European 14.46% 0.16%
Spanish es Indo-European 12.16% 0.13%
Vietnamese vi Austro-Asiatic 3.04% 0.08%
Indonesian id Austronesian 1.39% 0.03%
Japanese ja Japonic 0.00% 0.10%
Table 7: The language statistics of the pre-training corpora in BLOOM and LLaMA-2.
#Layer en zh fr es vi id ja
1 17 108 220 195 274 221 109
2 0 32 39 27 16 15 31
3 0 1 2 2 0 2 0
4 0 0 0 0 0 0 0
5 2 3 6 4 4 4 3
6 3 5 5 4 2 3 4
7 0 9 10 8 8 4 4
8 1 5 1 1 3 1 3
9 0 2 1 0 1 1 3
10 0 3 3 4 3 4 5
11 0 5 1 0 5 2 6
12 3 7 5 4 3 0 6
13 1 8 10 8 11 8 11
14 2 19 7 5 16 8 18
15 1 13 12 10 13 9 15
16 1 7 3 1 5 4 15
17 3 28 17 14 15 12 20
18 3 11 13 11 19 16 18
19 1 17 6 7 16 13 21
20 2 20 18 8 20 24 26
21 3 19 33 15 35 29 32
22 3 22 21 23 26 49 13
23 0 33 60 42 38 84 35
24 2 20 56 31 49 84 18
25 0 20 78 58 33 77 19
26 3 11 80 54 30 78 17
27 2 18 86 72 43 88 7
28 2 14 50 59 35 64 13
29 5 15 49 48 36 58 14
30 7 23 44 39 27 40 17
31 18 36 54 52 31 38 29
32 10 49 32 32 19 28 50
Table 8: Neuron number per layer of LLaMA-2 (7B).
#Layer en zh fr es vi id ja
1 60 127 222 189 248 184 206
2 9 162 177 118 187 69 305
3 0 2 1 2 1 2 1
4 0 1 1 1 0 2 2
5 0 3 0 1 0 0 3
6 1 3 2 2 3 3 5
7 0 5 1 1 4 3 6
8 3 7 3 3 3 1 9
9 2 18 7 7 10 3 9
10 0 12 9 6 8 5 9
11 0 15 18 17 11 8 12
12 2 5 3 2 5 3 9
13 1 7 2 1 3 2 9
14 0 5 3 2 4 0 10
15 1 7 3 3 7 5 8
16 3 25 20 14 22 11 31
17 3 30 16 11 28 21 32
18 4 40 40 31 35 25 47
19 1 26 23 13 22 19 44
20 1 24 14 16 14 9 35
21 1 28 19 17 22 17 34
22 3 43 26 13 40 37 55
23 3 32 23 10 22 24 49
24 1 28 20 12 20 32 27
25 1 24 29 8 23 25 29
26 3 27 40 32 27 42 31
27 2 40 63 41 31 64 36
28 4 20 50 43 21 48 30
29 2 25 78 48 19 71 26
30 0 25 89 88 38 75 19
31 2 21 72 52 46 77 13
32 3 16 83 60 36 80 14
33 0 23 69 55 31 61 19
34 1 27 47 54 35 69 16
35 1 20 69 58 41 67 23
36 1 21 60 54 42 58 27
37 5 22 33 32 31 47 11
38 14 40 58 52 48 44 36
39 8 57 43 35 26 27 51
40 15 105 47 51 38 44 97
Table 9: Neuron number per layer of LLaMA-2 (13B).
#Layer en zh fr es vi id ja
1 238 199 45 43 28 47 195
2 117 886 1056 1155 1589 897 1184
3 2 5 2 3 5 5 4
4 11 79 105 69 62 108 109
5 0 17 9 6 12 12 15
6 5 10 11 10 10 7 8
7 2 3 2 3 3 2 4
8 1 14 5 5 6 4 11
9 2 9 7 6 5 5 8
10 2 25 23 15 17 8 25
11 0 13 14 10 8 6 11
12 0 16 5 6 7 5 11
13 0 5 2 1 2 2 6
14 0 9 3 2 7 2 10
15 0 14 3 3 8 4 13
16 0 4 1 4 3 1 5
17 1 13 7 8 7 5 9
18 3 22 12 13 16 10 29
19 2 28 13 11 11 8 22
20 4 34 19 21 18 11 26
21 1 38 27 19 26 25 33
22 1 16 17 15 10 7 14
23 1 23 17 14 15 14 18
24 1 18 15 14 26 13 20
25 2 10 11 11 9 11 12
26 1 23 12 15 17 14 35
27 4 28 13 11 10 12 29
28 3 25 14 16 20 17 20
29 6 39 23 21 19 19 30
30 0 24 23 23 19 19 20
31 1 15 30 24 20 22 13
32 2 21 17 20 23 16 18
33 2 21 22 17 23 27 32
34 1 20 19 13 18 23 17
35 0 14 12 17 18 19 14
36 3 17 22 16 20 23 19
37 4 26 29 18 24 24 30
38 4 17 31 24 16 19 20
39 2 18 27 26 17 23 26
40 4 20 26 15 23 21 24
41 2 17 15 14 20 25 24
42 2 21 26 22 22 30 28
43 0 21 13 15 17 17 28
44 1 17 18 14 23 31 25
45 1 24 23 11 22 30 23
46 1 17 21 11 18 33 24
47 2 13 22 14 14 27 18
48 1 55 78 64 55 93 77
49 2 54 68 61 55 73 90
50 4 61 85 100 42 98 65
51 0 49 97 66 43 80 69
52 3 49 99 80 37 83 40
53 1 64 120 96 55 86 64
54 0 55 136 130 54 116 46
55 1 55 118 109 52 114 49
56 4 62 134 130 74 135 44
57 0 47 149 162 64 132 50
58 1 40 187 172 73 142 45
59 0 38 162 184 87 155 43
60 4 59 190 208 84 163 54
61 1 57 178 180 80 211 47
62 2 39 142 165 75 180 43
63 2 45 137 160 72 174 36
64 2 36 123 138 76 138 37
65 0 40 104 123 58 145 31
66 3 35 90 112 67 124 39
67 4 51 86 103 69 112 43
68 1 27 63 74 55 101 40
69 3 33 64 69 73 85 44
70 6 39 67 66 56 81 51
71 8 55 60 58 65 69 47
72 4 50 75 75 55 64 47
73 10 74 62 60 53 87 55
74 18 94 84 82 88 107 84
75 15 154 132 154 98 140 113
76 30 188 139 152 122 178 148
77 37 254 244 239 162 242 186
78 57 292 245 255 177 270 230
79 89 450 256 263 192 226 402
80 81 484 219 220 179 192 438
Table 10: Neuron number per layer of LLaMA-2 (70B).